I am having a difficult time figuring out part (c). Below is the question.

A monopolist’s inverse demand function is P = 150 – 3Q. The company produces output at two facilities; the marginal cost of producing at facility 1 is MC1(Q1) = 6Q1, and the marginal cost of producing at facility 2 is MC2(Q2) = 2Q2.

MR(Q) = 150- 6Q1 - 6Q2 Q1 = 5 Q2 = 15

c. Determine the profit-maximizing price.
Answer #1

Price P=150-3Q

when facility one produces 5 unit and 15unit from facility 2. Total production=5+15=20

Price=150-3(20)=90

Profit maximising price=90